OpenStudy (anonymous):

I will give medal!!! x over 2 = 15 over 6, for x

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (shaik0124):

is it x/2=15/6 ?

OpenStudy (anonymous):

yeah

OpenStudy (michele_laino):

\[\frac{ x }{ 2 }=\frac{ 15 }{ 6 }\] please multiply both sides by 2

OpenStudy (anonymous):

ok

OpenStudy (shaik0124):



OpenStudy (anonymous):

ok

OpenStudy (shaik0124):